# renew-dots,renew-matrix,nullify-dots

### 10 Dec 2023 v1

@@ Introduction
$\mathcal{T}^{(n)}_{ij}\!=\!\left(\!\!\left[\!\!\left(\left[\widehat{R}_{0}(n)\right]_{S}\right)^{-1}\right]_{ij}\!\!,\ldots,\!\!\left[\left(\left[\widehat{R}_{M}(n)\right]_{S}\right)^{-1}\right]_{ij}\!\right),\vspace{-5px}$
built from the inverse of the observed part of the lag moments. We have empirically observed that a certain normalized version of these features exhibits nontrivial separability and stability properties across a broad range of connectivity regimes and noise correlation levels $\beta$. Namely, coupled with the features $\mathcal{F}^{(n)}$, see equation (11), they aid in mitigating the oscillation and drift caused by the noise structure. More concretely, let $\mathcal{K}^{(n)}\overset{\Delta}{=}\mathcal{T}^{(n)}\times\mathcal{F}^{(n)}$ be the Cartesian product between the set of features defined in (Novel Features) and $\mathcal{F}^{(n)}$, i.e., $\mathcal{K}_{ij}^{(n)}=\left(\mathcal{F}_{ij}^{(n)},\mathcal{T}_{ij}^{(n)}\right)$ for each $ij$. Then, given Lemma $2$ in (Machado et al. 2023), the identifiability gap separating the features of connected pairs from features of disconnected pairs (refer to its formal definition in (Machado et al. 2023)) is larger for the features $\mathcal{K}^{(n)}$ than for either the features $\mathcal{T}^{(n)}$ or $\mathcal{F}^{(n)}$ w.h.p. under the sufficient condition in Theorem $2$. This tends to make the features $\mathcal{K}^{(n)}$ more amenable to generalization when using supervised methods to cluster them. Indeed, we have observed that the performance of supervised methods trained with a normalized version of $\mathcal{K}^{(n)}$ is better than the performance of the same method using either $\mathcal{T}^{(n)}$ or $\mathcal{F}^{(n)}$.
@@
We adopt the following training set ${\sf Tr}^{(n)}\overset{\Delta}{=}\left\{\left(\overline{\mathcal{K}}^{(n)}_{ij},\mathbf{1}_{\left\{A_{ij}\neq 0\right\}}\right)\right\}_{i\neq j}$, where we define the normalized features $\overline{\mathcal{K}}^{(n)}_{ij}:=\left[{\sf SC}\left(\mathcal{K}^{(n)}\right)\right]_{ij}$, with ${\sf SC}$ denoting the standard-scaler, i.e., the set $\mathcal{K}^{(n)}=\left\{\mathcal{K}^{(n)}_{ij}\right\}_{i\neq j}$ is centered and the entries of each feature $\mathcal{K}^{(n)}_{ij}$ are normalized to meet unit variance as described in (Pedregosa et al. 2011). The standard-scaler normalization helps to mitigate the drift of the features caused by the average noise cross-correlation $\beta$. That is, for training, we provide a normalized set of features $\overline{\mathcal{K}}^{(n)}$, with $\overline{\mathcal{K}}^{(n)}_{ij}$ being the feature of the pair $ij$ used as input to a Feed Forward Neural Network (FFNN) whose training output should be the ground-truth connectivity of the pair $ij$, i.e., $1_{\left\{A_{ij}\neq 0\right\}}$.
@@
Figure 3: Plots (a)-(f) contrast the accuracy of distinct methods as a function of the number of time series samples: i) (a)-(c) undirected networks (graph learning); (d)-(e) directed networks (causal inference); and (f) refers to a directed Brain connectome matrix. The hyperparameters $N$, $S$, $\beta$ and $p$ displayed at the top of the plots entail the underlying regime.
@@
We generate the time series data $\left\{\mathbf{y}(i)\right\}_{i=0}^{n}$ as done in (Machado et al. 2023), i.e., we first build a graph $G$ on $N$ nodes and then assign weights to the graph $G$ to obtain an $N\times N$ interaction matrix $A$ that is nonnegative and stable, i.e., $\rho(A)<1$. We create $G$ as a realization of an Erdős–Rényi random graph – without self-loops, i.e., $G_{ii}=0$ for all $i$ – with $N$ nodes and probability of edge or arrow placing $p$. The value of $p$ controls the sparsity of the underlying network. Then, the interaction matrix $A$ is characterized by: $A_{ij}=\alpha\frac{G_{ij}}{d_{\max}(G)}$, for $i\neq j$; and $A_{ii}=\rho-\sum_{k\neq i}A_{ik}$, for all $i$ where $d_{\max}(G)$ is the maximum inwards-degree of the underlying graph $G$ and $\alpha,\rho$ are stability hyperparameters with $10<\alpha\leq\rho<$ 1. This weight assignment is often referred to as Laplacian rule (Chung 1994). This generative process renders the networked dynamical system (1) stable and with a support graph of interactions given by $G$ – the main object of inference in our framework. To generate a directed graph $G$, we considered the realization of the Erdős–Rényi version for directed graphs also known as Binomial random graphs.
@@
We train FFNNs over one realization of a random graph model with $p=0.5$ and $N=50$ and apply this trained FFNN to networked systems with distinct connectivities obtained with different $p$. Further, for training, we generate datasets with distinct average noise cross correlation levels $\beta=0,5,10,\ldots,50$. Remarkably, our method generalizes to distinct connectivity regimes and different noise correlation levels as seen in the next Section.
@@
Fig. 3 demonstrates the overall performance superiority of our method in the colored regimens considered. Fig. 3 (f) refers to a human brain connectome, obtained via tractography (Škoch et al. 2022). In this latter case, we generate the time series $\left\{\mathbf{y}(n)\right\}_{n=1}^{M}$ for this real world network to infer its connectivity. The regimes (a)-(c) refer to undirected graphs obtained as realizations of an Erdős–Rényi random graph model, and (d),(e) refer to directed graphs, with the networks drawn from a Binomial random model.
The hyperparameters $N$, $S$, $\beta$ amd $p$ represent the number of nodes in the NDS, number of observed nodes, average cross correlation of the noise (average of the entries of ${\sf Osc}\left({\sf Off}(\Sigma_{x})\right)$), and the probability of edge or arrow placing in the corresponding Erdős–Rényi or Binomial random graph model (where applicable), respectively.
@@
We studied the problem of network identification from the times series data streaming from a linear networked dynamical system (NDS). We considered the challenging framework of partial observability under colored noise. We first devised a novel identifiability condition for linear NDS under this regime. Then, departing from the scalar based methods that are dominant in causal inference, we considered a feature based paradigm: to each pair of nodes, we assign a feature vector. We studied the impact of the noise structure on the identifiability properties of the features of (Machado et al. 2023) to propose novel features. We trained FFNNs with our features to cluster them and recover the network structure. The numerical experiments across distinct regimes of connectivity, observability and noise correlation demonstrate the competitive performance of our approach.
@@ Acknowledgments
The work of A. Santos was funded in part by FCT/MCTES through national funds and when applicable co-funded EU funds under the project UIDB/50008/2020 and project UIDP/00326/2020. This work is also partially supported by the U.S. National Science Foundation under Grant CCF 2327905. D. Rente and R. Seabra were funded in part by FCT - Foundation for Science and Technology, Portugal, under project UIDP/00326/2020. The source code for the numerical experiments can be found at https://github.com/seabrapt/brain˙underlying˙structure˙identification.

### 12 Feb 2024 v3

@@ Introduction
$\mathcal{T}^{(n)}_{ij}\!=\!\left(\!\!\left[\!\!\left(\left[\widehat{R}_{0}(n)\right]_{S}\right)^{-1}\right]_{ij}\!\!,\ldots,\!\!\left[\left(\left[\widehat{R}_{M}(n)\right]_{S}\right)^{-1}\right]_{ij}\!\right),$
built from the inverse of the observed part of the lag moments. We have empirically observed that a certain normalized version of these features exhibits nontrivial separability and stability properties across a broad range of connectivity regimes and noise correlation levels $\beta$. Namely, coupled with the features $\mathcal{F}^{(n)}$, see equation (11), they aid in mitigating the oscillation and drift caused by the noise structure. More concretely, define $\mathcal{K}_{ij}^{(n)}=\left(\mathcal{F}_{ij}^{(n)},\mathcal{T}_{ij}^{(n)}\right)$ for each $ij$. Then, given Lemma $2$ in (Machado et al. 2023), the identifiability gap separating the features of connected pairs from features of disconnected pairs (refer to its formal definition in (Machado et al. 2023)) is larger for the features $\mathcal{K}^{(n)}$ than for either the features $\mathcal{T}^{(n)}$ or $\mathcal{F}^{(n)}$ w.h.p. under the sufficient condition in Theorem $2$. This tends to make the features $\mathcal{K}^{(n)}$ more amenable to generalization when using supervised methods to cluster them.
@@
We adopt the following training set
${\sf Tr}^{(n)}\overset{\Delta}{=}\left\{\left(\overline{\mathcal{K}}^{(n)}_{ij},\mathbf{1}_{\left\{A_{ij}\neq 0\right\}}\right)\right\}_{i\neq j},\vspace{-5px}$ (16)
@@
where we define the normalized features $\overline{\mathcal{K}}^{(n)}_{ij}:=\left[{\sf SC}\left(\mathcal{K}^{(n)}\right)\right]_{ij}$, with ${\sf SC}$ denoting the standard-scaler, i.e., the set $\mathcal{K}^{(n)}=\left\{\mathcal{K}^{(n)}_{ij}\right\}_{i\neq j}$ is centered and the entries of each feature $\mathcal{K}^{(n)}_{ij}$ are normalized to meet unit variance as described in (Pedregosa et al. 2011). The standard-scaler normalization helps to mitigate the drift of the features caused by the average noise cross-correlation $\beta$. That is, for training, we provide a normalized set of features $\overline{\mathcal{K}}^{(n)}$, with $\overline{\mathcal{K}}^{(n)}_{ij}$ being the feature of the pair $ij$ used as input to a Feed Forward Neural Network (FFNN) whose training output should be the ground truth connectivity of the pair $ij$, i.e., $1_{\left\{A_{ij}\neq 0\right\}}$.
@@
Figure 3: Plots (a)-(f) contrast the accuracy of distinct methods as a function of the number of time series samples: $i)$ (a)-(c) undirected networks (graph learning); (d)-(e) directed networks (causal inference); and (f) refers to a directed Brain connectome matrix. The hyperparameters $N$, $S$, $\beta$ and $p$ displayed framed in the plots entail the underlying regime.
@@
We generate the time series data $\left\{\mathbf{y}(i)\right\}_{i=0}^{n}$ as done in (Machado et al. 2023), i.e., we first build a graph $G$ on $N$ nodes and then assign weights to the graph $G$ to obtain an $N\times N$ interaction matrix $A$ that is nonnegative and stable, i.e., $\rho(A)<1$. Namely, we create $G$ as a realization of an Erdős–Rényi random graph – without self-loops, i.e., $G_{ii}=0$ for all $i$ – with $N$ nodes and probability of edge or arrow placing $p$. The value of $p$ controls the sparsity of the underlying network. Then, the interaction matrix $A$ is characterized by: $A_{ij}=\alpha\frac{G_{ij}}{d_{\max}(G)}$, for $i\neq j$; and $A_{ii}=\rho-\sum_{k\neq i}A_{ik}$, for all $i$ where $d_{\max}(G)$ is the maximum inwards-degree of the underlying graph $G$ and $\alpha,\rho$ are stability hyperparameters with $10<\alpha\leq\rho<$ 1. This weight assignment is often referred to as Laplacian rule (Chung 1994). This generative process renders the NDS (1) stable and with a support graph of interactions given by $G$ – the main object of inference in our framework. To generate a directed graph $G$, we considered the realization of the Erdős–Rényi version for directed graphs also known as Binomial random graph model.
@@
We train FFNNs over one realization of a random graph model with $p=0.5$ and $N=50$ and apply this trained FFNN to networked systems with distinct connectivities obtained with different $p$. Further, for training, we generate datasets with distinct average noise cross-correlation levels $\beta=0,5,10,\ldots,50$. Remarkably, our method generalizes to distinct connectivity regimes and different noise correlation levels as seen in the next Section.
@@
Fig. 3 demonstrates the overall performance superiority of our method in the colored regimes considered. Fig. 3 (f) refers to a human brain connectome, obtained via tractography (Škoch et al. 2022). In this latter case, we generate the time series $\left\{\mathbf{y}(n)\right\}_{n=1}^{M}$ for this real world network to infer its connectivity. The regimes (a)-(c) refer to undirected graphs obtained as realizations of an Erdős–Rényi random graph model, and (d),(e) refer to directed graphs, with the networks drawn from a Binomial random graph model.
The hyperparameters $N$, $S$, $\beta$ and $p$ represent the number of nodes in the NDS, number of observed nodes, average cross-correlation of the noise (average of the off-diagonal entries of $\Sigma_{x}$), and the probability of edge or arrow placing in the corresponding Erdős–Rényi or Binomial random graph model (where applicable). Additional experiments for undirected graphs are in (Santos et al. 2024).
@@
We studied the problem of network identification from the times series data streaming from a linear networked dynamical system (NDS). We considered the challenging framework of partial observability under colored noise. We first devised a novel identifiability condition for linear NDS under this regime. Then, departing from the scalar based methods that are dominant in causal inference, we considered a feature based paradigm: to each pair of nodes, we assign a feature vector. We studied the impact of the noise structure on the identifiability properties of the features of (Machado et al. 2023) to propose novel features. We trained FFNNs with our features to cluster them and recover the causal network structure. The numerical experiments across distinct regimes of connectivity, observability and noise correlation demonstrate the competitive performance of our approach.
@@ Acknowledgments
This work is partially supported by the U.S. National Science Foundation under Grant CCF 2327905. The work of A. Santos was funded in part by FCT/MCTES through national funds and when applicable co-funded EU funds under the project UIDB/50008/2020 and project UIDP/00326/2020. José M. F. Moura was partially supported by the U.S. National Science Foundation under Grant CCN 1513936. D. Rente and R. Seabra were funded in part by FCT - Foundation for Science and Technology, Portugal, under project UIDP/00326/2020. The source code for the numerical experiments can be found at https://github.com/seabrapt/brain˙underlying˙structure˙identification.